OpenStudy (anonymous):

can anyone help me with this? Find the general term of the set. {8, 12, 16, 20, 24, . . .}

OpenStudy (anonymous):

4

OpenStudy (anonymous):

12-8=? 16-12=? 20-16=? ...........

OpenStudy (anonymous):

I think it would be 4n+4... but I am unsure

jimthompson5910 (jim_thompson5910):

you are correct, the nth term is 4n+4 (where n starts at n=1)

OpenStudy (anonymous):

8n 4n + 4 3n + 5 5n + 3 out of these options

OpenStudy (anonymous):

ok thank you so much, I hate math and i am taking an online class, my instructor seems to jump around the book so much as oppose to teaching one lesson at a time.

OpenStudy (anonymous):

4(1)+4=? 4(2)+4=? 4(3)+4=?

OpenStudy (anonymous):

online is pretty tough gdluck

OpenStudy (anonymous):

thanks... and one more: Let U = {5, 10, 15, 20, 25, 30, 35, 40} A = {5, 10, 15, 20} B = {25, 30, 35, 40} C = {10, 20, 30, 40}. Find A ⋂ B.

jimthompson5910 (jim_thompson5910):

A ⋂ B = set of all things that are in BOTH set A and B

OpenStudy (anonymous):

I have never seen the upside down "u" symbol before.. ahhh thanks

jimthompson5910 (jim_thompson5910):

it's known as the intersection symbol

OpenStudy (anonymous):

thank you i really appreciate this!

OpenStudy (anonymous):

Jim, can you help with another one? I am pretty sure I figured it out but just to make sure... Which sentence is not a statement? 2 x 4 = 7 8 + 2 Jan and Ella are neighbors. Bill is a photographer

OpenStudy (anonymous):

I said it was 8+2

jimthompson5910 (jim_thompson5910):

8 + 2 is just an expression not a statement Note: 2 x 4 = 7 is a statement even though it's false

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

a statement is something that is either true or false

OpenStudy (anonymous):

I figured that it was because there wasnt an actual answer, false or not

jimthompson5910 (jim_thompson5910):

but 8+2 is neither true or false

OpenStudy (anonymous):

that's what I meant, because there wasn't an answer to make it either true or false

OpenStudy (anonymous):

Select a counterexample to show that the statement is false. When an odd number is added to the product of two even numbers, the result will be even. 4•5 + 1 = 21. 21 is not even. 4•2 + 3 = 11. 11 is not even. 4•5 + 2 = 22. 22 is even. 3•5 + 4 = 19.19 is not even. on this one I chose the second one, is that accurate?

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (anonymous):

I am also doing sets and subsets, can I give you another one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Which statement is false? (Points : 1) a ∈ {a, b, c} Ø ⊂ {a, b, c} Ø ⊆ {a, b, c} Ø ∈ {a, b, c} I chose the last one, but I my book

OpenStudy (anonymous):

does not seem to explain the set and subset stuff clearly

jimthompson5910 (jim_thompson5910):

a ∈ {a,b,c} means "the element 'a' is in the set {a,b,c}" and this is true

OpenStudy (anonymous):

I got that one, but which one is false? I don't really know what the other ones mean

jimthompson5910 (jim_thompson5910):

the ⊂ means "is a proper subset of"

jimthompson5910 (jim_thompson5910):

the ⊆ means "is a subset of"

OpenStudy (anonymous):

what does \[\emptyset\] mean?

jimthompson5910 (jim_thompson5910):

that's the empty set

jimthompson5910 (jim_thompson5910):

since the empty set is a subset of every set, this means that choices 2 and 3 are also true

jimthompson5910 (jim_thompson5910):

but choice 4 is not true because the empty set isn't listed in the set on the right side

OpenStudy (anonymous):

ok, awesome! so I was accurate when I chose the last one?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Which of the following phrases does not describe the number -¾? (Points : 1) rational number irrational number real number negative number I chose irrational on this one,

jimthompson5910 (jim_thompson5910):

good, it's not irrational since it's rational

OpenStudy (anonymous):

how are you at explaining how to solve word problems?

jimthompson5910 (jim_thompson5910):

show me the problem and we'll find out, it depends on the problem

OpenStudy (anonymous):

Upon examining the contents of 38 backpacks, it was found that 23 contained a black pen, 27 contained a blue pen, and 21 contained a pencil, 15 contained both a black pen and a blue pen, 12 contained both a black pen and a pencil, 18 contained both a blue pen and a pencil, and 10 contained all three items. How many backpacks contained exactly two of the three writing instruments? (Points : 1) 11 2 3 15 I am trying to write it out so it makes sense

jimthompson5910 (jim_thompson5910):

Are you familiar with venn diagrams?

OpenStudy (anonymous):

yeah and i tried to work out a problem from my book and eventually figured it out, but let me know if this is how I would start. My instructor gave us a study guide and i am trying to replace the figures from that one with these.

OpenStudy (anonymous):

woud I start with this? \[15-10=5\] \[12-10=2\] \[18-10=8\]

jimthompson5910 (jim_thompson5910):

and those figures refer to what

jimthompson5910 (jim_thompson5910):

eg: 5 refers to number of backpacks with black and blue pens only

OpenStudy (anonymous):

the 15 is for black and blue pens, the 12 is for black and pencil, and the 18 is a blue pen and pencil

jimthompson5910 (jim_thompson5910):

ok, so you would use that to fill in the diagram. These new numbers would go in between pairs of circles

OpenStudy (anonymous):

subtracting the 10 because we are trying to find out how many have 2 items not three, so we take them out of the equation. then we should just add up 5,2,and 8 and get 15

jimthompson5910 (jim_thompson5910):

so 5 black and blue only, 2 black and pencil only, and 8 blue and pencil

OpenStudy (anonymous):

is this a two or three circle diagram?

jimthompson5910 (jim_thompson5910):

3 overlapping circles

jimthompson5910 (jim_thompson5910):

now we need to find out how many backpacks there are with only one item, how do we do that

OpenStudy (anonymous):

would it be: \[23-10-5-2=7 for black pens only\]

jimthompson5910 (jim_thompson5910):

good, there are 7 backpacks with black pens only

jimthompson5910 (jim_thompson5910):

do the same for the other two items

OpenStudy (anonymous):

ok, phew, I actually did the diagram right

OpenStudy (anonymous):

so the question asks for the back packs with 2 items, so that is 15 right

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

ok so for the other ones it would be 1 with pencil, and 4 with blue pens

jimthompson5910 (jim_thompson5910):

oh wait I made a typo, there should be only 6 backpacks with black pens only

OpenStudy (anonymous):

oh yeah, that was my fault

jimthompson5910 (jim_thompson5910):

1 pencil, 4 blue, so your other numbers are correct

OpenStudy (anonymous):

another question uses the same variables, but wants to know, . How many backpacks contained none of the three writing instruments?

jimthompson5910 (jim_thompson5910):

ok you have the number of backpacks which have all 3, two items, or just one item right?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

add up all of those individual numbers

jimthompson5910 (jim_thompson5910):

what do you get when you do that?

OpenStudy (anonymous):

37

jimthompson5910 (jim_thompson5910):

I'm getting 36, what are your numbers

jimthompson5910 (jim_thompson5910):

One item only 6, 1, 4 Two items 5, 2, 8 All three 10

OpenStudy (anonymous):

oh I found my mistake

OpenStudy (anonymous):

6 black pens and not 7

jimthompson5910 (jim_thompson5910):

oh ok

jimthompson5910 (jim_thompson5910):

so there are 36 items that have either a) one item only b) two items, or c) all three items So subtract this from the total 38 to get 38-36 = 2 So there are two backpacks that have none of these items

OpenStudy (anonymous):

ok, I got it! it is easier when you do the diagram

jimthompson5910 (jim_thompson5910):

yeah it helps you see where everything goes

OpenStudy (anonymous):

I have a question about subsets again.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

. Find the number of subsets the set has. {1, 2, 3, 4, 5, 6, 7, 8, 9}

OpenStudy (anonymous):

what is the easiest way to figure this out?

jimthompson5910 (jim_thompson5910):

If a set has 'n' elements, then the set will have \[2^n\] subsets

jimthompson5910 (jim_thompson5910):

In this case, n=9, so there are \[2^9 = 512\] subsets

OpenStudy (anonymous):

and it is always to the power of the number of elements?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

well 2 to that power

OpenStudy (anonymous):

is it 2 because every set is a subset of itself?

jimthompson5910 (jim_thompson5910):

it's 2 because each element is either in a subset or it's not in a subset (ie either a yes or a no)

OpenStudy (anonymous):

ok...this is another one that confuses me: Which set is infinite? {x | x ∈ N and x < 20} {x | x is an improper fraction} {3, 5, 7, . . ., 99} Ø

jimthompson5910 (jim_thompson5910):

{3, 5, 7, . . ., 99} is finite because it starts with 3 and counts by 2 until it hits 99

jimthompson5910 (jim_thompson5910):

{x | x ∈ N and x < 20} is finite because this is the set of positive numbers from 1 to 19

OpenStudy (anonymous):

i ruled that one out because of just that

jimthompson5910 (jim_thompson5910):

Ø is finite because there are no elements in this set (and it's not infinite)

OpenStudy (anonymous):

I would think it is the one about the improper fraction

jimthompson5910 (jim_thompson5910):

{x | x is an improper fraction} is infinite because there are an infinite number of improper fractions

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Which property of real numbers does the following equation demonstrate? 4(x – 7) = 4x – 28 I said this is distributive property